OpenStudy (anonymous):

Solve: x - 3 = 5x + 13 My answer is -6 Am I right? :)

OpenStudy (ja1):

-4 X - 3=5X+13 X = 5X + 16 -4X=16 divide it all by -4 -4X divided by -4 is X 16 divided by -4 is -4 X=-4 Answer is -4

OpenStudy (anonymous):

Ok thank you! :)

OpenStudy (ja1):

no prob :)